OpenStudy (anonymous):

Simplify: [(-5^-1)(2^-2)]^2

OpenStudy (anonymous):

A negative power is the same as 1/base so: [(1/-5)(1/2)]^2 = (1/-10)^2 = 1/100

OpenStudy (anonymous):

i got .oo25 or 1/400

jhonyy9 (jhonyy9):

((1/-5)(1/4))*2 = (1/-20)*2 = 1/400